# We’ll Fix it in Post: Improving Text-to-Video Generation with Neuro-Symbolic Feedback

### v2

@@ Abstract
Current text-to-video (T2V) generation models are increasingly popular due to their ability to produce coherent videos from textual prompts. However, these models often struggle to generate semantically and temporally consistent videos when dealing with longer, more complex prompts involving multiple objects or sequential events. Additionally, the high computational costs associated with training or fine-tuning make direct improvements impractical. To overcome these limitations, we introduce NeuS-E, a novel zero-training video refinement pipeline that leverages neuro-symbolic feedback to automatically enhance video generation, achieving superior alignment with the prompts. Our approach first derives the neuro-symbolic feedback by analyzing a formal video representation and pinpoints semantically inconsistent events, objects, and their corresponding frames. This feedback then guides targeted edits to the original video. Extensive empirical evaluations on both open-source and proprietary T2V models demonstrate that NeuS-E significantly enhances temporal and logical alignment across diverse prompts by almost $40\%$.
Figure 1: The teaser demonstrates that the original video generation does not adhere to the prompt temporally. However, after editing the video using NeuS-E, it aligns with the prompt while preserving well-generated portions. The color of the frame border corresponds to the color of the events and semantics in the text prompt. In the first example at the top, the original video omits the scene where a cyclist reaches a park and enjoys the scenery. However, after editing with NeuS-E, the park becomes visible, and the cyclist is shown arriving. In the second example at the bottom, the sun remains visible in the original generation. However, NeuS-E successfully edits it, causing the sun to set and disappear behind the mountain.
footnotetext: *Equal contribution (Order determined by coin toss).
Corresponding author: minkyu.choi@utexas.edu
@@ 1 Introduction
Consider attempting to generate the following complex scenario from a text-to-video (T2V) model:
“An autonomous vehicle navigates an intersection, waits for a pedestrian to cross, and then proceeds.”
@@
However, current T2V models fail catastrophically when prompted to generate events in a temporal order. Even after extensive re-prompting, these models fail to generate well-aligned videos (see Fig. 1). Directly fine-tuning these models to improve temporal alignment is computationally expensive and often fails to generalize across different prompts and temporal event relationships. What if, instead of fine-tuning T2V models, we formally identify the problematic video segments and surgically refine only those parts with targeted edits?
Research on T2V generation exists, but existing attempts mainly focus on semantic improvements. These methods modify rewards in diffusion models, tweak attention maps, and perform low-level architectural modifications. Though they improve visual quality, they cannot fix temporal misalignment or missing events. Additionally, such modifications are infeasible without direct access to the model weights and training code, which is often unavailable for closed-source models such as Gen3 [21] and Pika [68].
Likewise, methods that rely on explicit re-prompting to optimize for video evaluation metrics do not work well for generating temporally aligned videos. Metrics such as VBench [34], EvalCrafter [54], and T2V-Bench [39] focus primarily on aesthetic and spatial fidelity without direct evaluation of temporal correctness. To solve this problem, NeuS-V [72] introduced a metric for temporal fidelity using neuro-symbolic formal verification methods, offering a more rigorous way to evaluate temporal consistency in synthetic videos. This leads to the following question:
“Can we improve the temporal fidelity of T2V generation using neuro-symbolic feedback?”
In this paper, we propose NeuS-E, a zero-training method to pinpoint weak parts of a generated video and refine them selectively through structured neuro-symbolic feedback, achieving stronger alignment without retraining generation models. Our key insight is to take NeuS-V beyond evaluation and turn it into an active refinement tool. Instead of treating generated videos as static outputs, we introduce an iterative feedback-driven editing loop. First, NeuS-E identifies weakly satisfied events or semantics in the generated video, and their corresponding segments are iteratively refined while ensuring smooth transitions and improved temporal consistency. Through this, we obtain a video that better aligns with the temporal specifications in the prompt without requiring endless re-prompting for regeneration. Our key contributions are as follows:
@@
  • We introduce NeuS-E which automatically corrects weak segments through neuro-symbolic feedback by identifying and editing the keyframe, then regenerating misaligned video segments to improve temporal alignment.
  • We demonstrate that NeuS-E significantly boosts the temporal fidelity of both open-source (CogVideoX) and closed-source (Gen-3, Pika) models—all with zero additional training.
@@ 2 Related Work
Following the recent successes in text-to-image generation, text-to-video models such as SORA [67] from OpenAI, GEN-3 Alpha [70] from Runway, Kling [48] from Kling AI, Veo [73] from Google, Wan [76], and PIKA [49] have seen significant advancements. While closed-source and inaccessible for the community to conduct research, these models aim to generate coherent videos from textual prompts by leveraging diffusion-based architectures [29] or autoregressive approaches like the CogVideo series [31, 87]. For instance, Phenaki [78], I2VGen-XL [95], and Esser et al. [18] were some of the earlier works that focus on efficient latent-space based video synthesis based on text prompts. Currently, text-to-video models such as the VideoCrafter series [7, 8], Make-your-Video [82], Moonshot [92], Show-1 [93], VDM [30], LVDM [27], HunyuanVideo [44], and VideoFactory [80] learn both the spatial and temporal modules. Other methods such as Align Your Latent [3], AnimateDiff [25], and Text-to-Video [42] zero only train a temporal adapter or module with a fixed spatial module. For a more detailed survey we refer you to [12]. In contrast to these methods, NeuS-E proposes a training-free approach to video generation, for temporally complex prompts.
@@
Many recent works have proposed training-free frameworks to improve or stylize text-to-video generation. These methods, such as [37, 91, 42, 85, 79, 24, 23], leverage text-to-image (T2I) models [63, 4, 94] to edit video frames. Approaches like TokenFlow [22], Rerender [85], and VideoP2P [52] aim to improve temporal coherence by identifying key frames [6] or rectifying cross-frame attention maps and features [37, 69]. Other methods refine video generation by improving text prompts [43, 57], combining text and images for edits [92], or employing video diffusion models like Dreamix [64] to enhance consistency. Despite these advancements, most editing-based methods focus solely on improving the temporal fidelity of objects or semantic features within a video. They fail to address the temporal coherence across multiple events generated in the video. For example, these methods do not ensure alignment between event sequences (e.g., “a person enters a car before driving away”). To bridge this gap, we propose NeuS-E, which leverages neuro-symbolic feedback to augment prompts and guides edits for generating synthetic videos that satisfy complex temporal relationships across multiple coherent events.
@@
Existing approaches often rely on distribution-based metrics from features such as FID, FVD, and CLIPSIM [74, 53, 36, 5, 32, 90], but these struggle with prompts requiring complex reasoning or logical consistency. Recent works [45, 55, 96, 50] leverage Large Language Models (LLMs) to decompose prompts into Visual Question Answering (VQA) pairs scored by Vision-Language Models (VLMs). Other methods incorporate vision transformers and temporal networks [33, 39, 20, 14], or ensembles of visual quality metrics as seen in EvalCrafter [54], FETV [56], and T2V-Bench [20]. Among these, VBench [34] is emerging as a leading benchmark by evaluating across diverse dimensions. While VBench does assess temporal quality it does so primarily from an aesthetic perspective, focusing on aspects like consistency, flickering, and movement. In contrast, NeuS-E utilizes a recent work NeusV [72] that addresses these gaps by formally verifying temporal fidelity using Temporal Logic (TL) specifications over automaton representations of synthetic videos to better evaluate T2V alignment.
@@
Formal verification methods that construct symbolic representation of events, tasks and semantics have been explored across various domains, including video classification [19, 77], event detection [60, 84, 51], video question-answering [88, 10], robotics [75, 26, 46], and autonomous driving [38, 61]. These approaches employ diverse techniques such as constructing graph-based structures for relational reasoning [89, 59, 83], utilizing latent-space representations as symbolic abstractions [71, 2, 47], or leveraging formal language frameworks [1] to define precise specifications. In contrast to these methods, NeuS-E aims to leverage formal verification to provide structured feedback by identifying text-based edit instructions and key-frames, to improve T2V generation without any training.
@@ 3 Preliminaries
We begin with a comprehensive running example of generating a promotional video using the prompt: ‘A person meditates by the lake and, a few seconds later, stands up for a moment before leaving’ to describe the following terminologies.
@@
Temporal Logic (TL) is an expressive formal language that combines logical and temporal operators to express time-dependent statements [17, 58]. A TL specification or formula is structured around three key components: ➊ a set of atomic propositions, ➋ first-order logic operators, and ➌ temporal operators. Atomic propositions are fundamental statements that evaluate to either True or False and serve as building blocks for more complex expressions. First-order logic operators include AND ($\wedge$), OR ($\vee$), NOT ($\neg$), IMPLY ($\Rightarrow$)., and the temporal operators consist of ALWAYS ($\Box$), EVENTUALLY ($\diamondsuit$), NEXT ($\mathsf{X}$), UNTIL ($\mathsf{U}$), etc. The set of atomic propositions $\mathcal{P}$, and the TL specification $\Phi$ of our running example are as follows:
$\begin{array}[]{rl}\mathcal{P}&=\{\text{person is meditating},\ \text{lake shore},\\ &\text{person is meditating},\text{person is walking away}\}.\\ \Phi&=(\text{person is meditating}\wedge\text{lake shore})\ \mathsf{X}\\ &\text{person is standing}\ \mathsf{X}\ \text{person is walking away}.\end{array}$ (1)
@@
A video automaton formally represents a video sequence as a sequence of states and transitions [13, 72, 86]. It is a form of Discrete-Time Markov Chain (DTMC) [65, 41] used to model generated videos as each frame’s transition depends only on the current frame. Since video sequences are inherently discrete, finite, and temporal, DTMCs can approximate frame transitions. To this end, we define the set of states in the video as $Q$ and this frame transition function as $\delta:Q\times Q\to[0,1]$ where, given two states $q_{1}\in Q$ and $q_{2}\in Q$, $\delta(q_{1},q_{2})\in[0,1]$ gives the probability of transitioning from $q_{1}$ to $q_{2}$. Now, the video automaton, $\mathcal{A}_{\mathcal{V}}$, is the tuple $\mathcal{A}_{\mathcal{V}}=(Q,q_{0},\delta,\lambda)$ where the initial state is $q_{0}\in Q$ and the label function is $\lambda:Q\to 2^{|\mathcal{P}]}$.
@@
Formal verification ensures formal guarantees that the system satisfies the desired specification. [15, 35]. It necessitates a formal representation of the system, such as a finite-state automaton (FSA). Given the video automaton $\mathcal{A}_{\mathcal{V}}$, a path in a video is defined as a sequence of states $q_{0}q_{1}(q_{2})^{\omega}$ starting from the initial state $q_{0}$ and $\omega$ denotes repetition. A trace corresponds to the sequence of labels $\lambda(q_{0})\lambda(q_{1})\lambda(q_{2})\cdots\in(2^{|\mathcal{P}|})^{\omega}$ associated with the states along a path, where $\lambda(q)$ represents the labeling function that maps each state $q$ to a subset of atomic propositions from the set $\mathcal{P}$. The trace captures the progression of events or properties over time as observed along the path. Next, we apply probabilistic model checking [1] to determine how much the trace starting from the initial state satisfies the temporal logic (TL) specification $\Phi$. Using these formal representations, we evaluate the videos.
@@
Text prompts are converted to temporal logic conversion using LLMs [9, 16, 62, 86, 72] for subsequent analysis. We denote the LLM-based text-to-temporal logic (T2TL) function as $f_{\text{T2TL}}:T\to(\mathcal{P},\Phi)$ to decompose a text prompt $T$ into a TL specification $\Phi$ and a set of propositions $\mathcal{P}$. A prompt example is presented in the Appendix.
@@
Definition 1 (Video Automaton Construction).
Given a generated video $\mathcal{V}$ and a set of atomic propositions $\mathcal{P}$, we construct a video automaton. For each proposition $p_{i}\in\mathcal{P}$ and frame $\mathcal{F}_{n}\in\mathcal{V}$, a VLM computes a semantic confidence score: $\mathcal{M}_{\text{VLM}}:p_{i}\times\mathcal{F}_{n}\to c_{i,n}$, where $c_{i,n}\in[0,1]$ represents the confidence of $p_{i}$ in $\mathcal{F}_{n}$. For each frame $n$, we define the confidence set as: =$\mathbb{C}=\{c_{i,n}\mid p_{i}\in\mathcal{P},\mathcal{F}_{n}\in\mathcal{V}\ \}$. The video automaton $\mathcal{A}_{\mathcal{V}}$ is then generated by applying a function $\xi$ that processes the set of propositions $\mathcal{P}$ along with the confidence scores across all frames:
@@
Definition 2 (Satisfaction Probability).
Given a video automaton $\mathcal{A}_{\mathcal{V}}$ and a temporal logic specification $\Phi$, the satisfaction probability $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$ is computed by verifying $\mathcal{A}_{\mathcal{V}}$ against $\Phi$ using STORM [28, 40], which uses probabilistic computation tree logic (PCTL). This verification process is formalized as a probabilistic model checking function:

### v4

# We’ll Fix it in Post: Improving Text-to-Video Generation with Zero Training

@@ Abstract
Current text-to-video (T2V) generation models are increasingly popular due to their ability to produce coherent videos from textual prompts. However, these models often struggle to generate semantically and temporally consistent videos when dealing with longer, more complex prompts involving multiple objects or sequential events. Additionally, the high computational costs associated with training or fine-tuning make direct improvements impractical. To overcome these limitations, we introduce NeuS-E, a novel zero-training video refinement pipeline that leverages neuro-symbolic feedback to automatically enhance video generation, achieving superior alignment with the prompts. Our approach first derives the neuro-symbolic feedback by analyzing a formal video representation and pinpoints semantically inconsistent events, objects, and their corresponding failure frames with respect to the prompt. This feedback then guides targeted edits to the original video. Extensive empirical evaluations on both open-source and proprietary T2V models demonstrate that NeuS-E significantly enhances temporal and logical alignment across diverse prompts by almost $40\%$.
@@ 1 Introduction
Imagine generating the following complex scenario from a text-to-video (T2V) model:
  • “An autonomous vehicle crosses an intersection after waiting for a pedestrian to cross."
@@
Current research in T2V generation has primarily focused on the first two aspects. Existing methods improve visual quality and semantic accuracy by modifying rewards in diffusion models, tweaking attention maps, or making low-level architectural changes. Although these enhance visual fidelity, they fail to fix temporal misalignments and are often infeasible for models like Gen3 [gen3] and Pika [pika], which lack accessible model weights.
Figure 1: NeuS-E improves the text-to-video (T2V) temporal alignment. The border color of the frames corresponds to the identified events. Vanilla T2V models (top video) fail to generate the sunset behind the mountains. NeuS-E systematically identifies and surgically corrects this video segment to improve the temporal fidelity of the synthetic video with targeted feedback.
However, state-of-the-art T2V models fail catastrophically when prompted to generate events in a specific temporal order. Training-free methods that utilize excessive remarking to improve traditional T2V evaluation benchmarks, such as VBench [huang2024vbench], fail to improve temporal alignment since these metrics prioritize visual aesthetics. In contrast, NeuS-V [sp_2025_CVPR] introduces a neuro-symbolic method that rigorously quantifies a text-to-video model’s temporal alignment with respect to the prompt. Taking inspiration from this work, we pose the following research question:
  • “Can we leverage neuro-symbolic feedback to surgically refine temporally misaligned video segments through targeted edits, thereby improving text-to-video temporal alignment?"
We introduce NeuS-E, a zero-training framework for targeted video refinement. Inspired by NeuS-V, we devise a neuro-symbolic feedback loop that disentangles atomic events and objects (termed propositions) with their prompted temporal order (termed specifications). Therefore, instead of treating videos as static outputs, NeuS-E identifies video segments that weakly satisfy the objects and events in the prompt, edits the corresponding keyframes, and regenerates only these misaligned portions to improve temporal alignment with respect to the prompt. This iterative process produces a video that aligns with the prompt’s temporal requirements without retraining the generative model.
A key insight behind our approach is that improving T2V generation does not necessarily require modifying the generative model itself. Instead, we treat generation errors as diagnosable failures of a temporal specification. By formally verifying the generated video against a temporal logic representation of the prompt, we can identify which proposition is violated and where the failure occurs in time. This diagnostic signal allows us to target only the portion of the video responsible for the violation, enabling localized refinement rather than repeated full-video regeneration. Our key contributions are summarized as follows:
@@
  • We introduce NeuS-E, which automatically corrects weak segments through neuro-symbolic feedback by identifying and editing the keyframe, then regenerating misaligned video segments to improve temporal alignment.
  • We demonstrate that NeuS-E significantly boosts the temporal fidelity of both open-source (CogVideoX) and closed-source (Gen-3, Pika) models, all with zero additional training.
@@ 2 Related Work
Text-to-video models such as SORA [openai2024sora], GEN-3 Alpha [runway2024gen3], Kling [kling], Veo [veo2024], Wan [wan2.1], and PIKA [pikaai2024] have seen significant advancements. The underlying research follows two main training strategies. The first involves end-to-end training of spatial and temporal modules using diffusion [ho2020denoising] or autoregressive [hong2022cogvideo, yang2024cogvideox] architectures, encompassing early latent-space works [villegas2022phenaki, zhang2023i2vgen, esser2023structure] and numerous modern systems [chen2023videocrafter1opendiffusionmodels, chen2024videocrafter2, xing2024make, zhang2024moonshot, zhang2024show, Ho2022VideoDM, he2023latentvideodiffusionmodels, kong2025hunyuanvideosystematicframeworklarge, wang2024videofactory, wan2.1]. The second strategy efficiently adapts pre-trained image models by training only a lightweight temporal module [blattmann2023align, Guo2023AnimateDiffAY, khachatryan2023text2videozerotexttoimagediffusionmodels]. For a broader survey, see [cho2024sora]. In contrast to these methods, NeuS-E proposes a training-free approach to video generation, for temporally complex prompts.
@@
Training-free frameworks improve text-to-video generation by leveraging text-to-image (T2I) models [Meng2021SDEditGI, Brooks2022InstructPix2PixLT, zhang2023adding] to edit video frames [Jeong2023GroundAVideoZG, Zhang2024MoonshotTC, khachatryan2023text2videozerotexttoimagediffusionmodels, Yang2023RerenderAV, wang2024zeroshotvideoeditingusing, goel2024improving, goel2024syndiff]. Most of these methods focus on improving the temporal coherence of individual objects by rectifying cross-frame attention maps and features [Geyer2023TokenFlowCD, Jeong2023GroundAVideoZG, Qi2023FateZeroFA, Yang2023RerenderAV, Liu2023VideoP2PVE], by refining text prompts [kim2024free2guidegradientfreepathintegral, luo2025enhanceavideobettergeneratedvideo], by editing keyframes [Ceylan2023Pix2VideoVE, zhang2024moonshot], or employing video-to-video diffusion models [molad2023dreamixvideodiffusionmodels]. While effective for object-level consistency, these approaches fail to enforce the logical order of multiple, distinct events (e.g., ensuring a person enters a car before driving away). To bridge this critical gap, NeuS-E uses neuro-symbolic feedback to guide edits, ensuring the final video satisfies complex temporal relationships across the entire event sequence.
@@
Current Text-to-Video (T2V) evaluation methods include distribution-based metrics (FID, FVD, CLIPSIM) [shin2024lostmelodyempiricalobservations, liu2023fetvbenchmarkfinegrainedevaluation, jain2024peekaboointeractivevideogeneration, bugliarello2023storybenchmultifacetedbenchmarkcontinuous, hu2022makemovecontrollableimagetovideo, yu2023celebvtextlargescalefacialtextvideo], LLM-based VQA scoring [kou2024subjectivealigneddatasetmetrictexttovideo, liu2024evalcrafterbenchmarkingevaluatinglarge, zhang2024benchmarkingaigcvideoquality, Li_2024_CVPR], and metric ensembles for visual and spatial quality like VBench, EvalCrafter, FETV, and T2V-Bench [huang2024vbench, liu2024evalcrafter, liu2024fetv, ji2024t2vbench, feng2024tc, huang2023t2i, chu2024sora]. While comprehensive, these approaches primarily assess aesthetic temporal quality, such as motion and flickering [huang2024vbench], rather than logical event sequencing. In contrast, NeuS-E leverages NeuS-V [sp_2025_CVPR] to address this gap, using Temporal Logic (TL) to formally verify complex temporal relationships.
@@
Formal verification methods are used to construct symbolic representations of events and tasks across various domains. These methods have been applied in video understanding [short1, short2, event-detection-video-stream, cnn-event-detect, lstm-event-detect, neural-symbolic, chen2022comphy], robotics [shoukry2017linear, hasanbeig2019reinforcement, kress2009temporal], and autonomous driving [jha2018safe, mehdipour2023formal], using techniques like graph-based reasoning [graphical-model-relationship-detection, visual-symbolic, long3], latent-space abstractions [symbolic-high-speed-video, BertasiusWT21, neural-symbolic-cv], or formal languages [Baier2008]. In contrast, NeuS-E leverages formal verification to provide structured feedback—identifying text-based edit instructions and keyframes—to improve Text-to-Video generation without any training.
@@ 3 Preliminaries
We begin with a comprehensive running example of generating a promotional video using the prompt: “A person meditates by the lake and, a few seconds later, stands up for a moment before leaving" to describe the following terminologies.
@@
Temporal Logic (TL) is an expressive formal language that combines logical and temporal operators to express time-dependent statements [Temporal-and-Modal-Logic, Manna]. A TL specification or formula is structured around three key components: ➊ a set of atomic propositions, ➋ first-order logic operators, and ➌ temporal operators. Atomic propositions are fundamental statements that evaluate to either True or False and serve as building blocks for more complex expressions. First-order logic operators include AND ($\wedge$), OR ($\vee$), NOT ($\neg$), IMPLY ($\Rightarrow$)., and the temporal operators consist of ALWAYS ($\Box$), EVENTUALLY ($\diamondsuit$), NEXT ($\mathsf{X}$), UNTIL ($\mathsf{U}$), etc. The set of atomic propositions $\mathcal{P}$, and TL specification $\Phi$ of our running example are:
[…]
@@
A video automaton formally represents a video sequence as a sequence of states and transitions [choi2025towards, sp_2025_CVPR, yang2023specification]. It is a form of Discrete-Time Markov Chain (DTMC) [norris1998markov, kemeny1960finite] used to model generated videos as each frame’s transition depends only on the current frame. Since video sequences are inherently discrete, finite, and temporal, DTMCs can approximate frame transitions. To this end, we define the set of states in the video as $Q$ and this frame transition function as $\delta:Q\times Q\to[0,1]$ where, given two states $q_{1}\in Q$ and $q_{2}\in Q$, $\delta(q_{1},q_{2})\in[0,1]$ gives the probability of transitioning from $q_{1}$ to $q_{2}$. Now, the video automaton, $\mathcal{A}_{\mathcal{V}}$, is the tuple $\mathcal{A}_{\mathcal{V}}=(Q,q_{0},\delta,\lambda)$ where the initial state is $q_{0}\in Q$ and the label function is $\lambda:Q\to 2^{|\mathcal{P}]}$.
@@
Formal verification ensures formal guarantees that the system satisfies the desired specification. [clarke1999model, huth2004logic]. It necessitates a formal representation of the system, such as a finite-state automaton (FSA). Given the video automaton $\mathcal{A}_{\mathcal{V}}$, a path in a video is defined as a sequence of states $q_{0}q_{1}(q_{2})^{\omega}$ starting from the initial state $q_{0}$ and $\omega$ denotes repetition. A trace corresponds to the sequence of labels $\lambda(q_{0})\lambda(q_{1})\lambda(q_{2})\cdots\in(2^{|\mathcal{P}|})^{\omega}$ associated with the states along a path, where $\lambda(q)$ represents the labeling function that maps each state $q$ to a subset of atomic propositions from the set $\mathcal{P}$. The trace captures the progression of events or properties over time as observed along the path. Next, we apply probabilistic model checking [Baier2008] to determine how much the trace starting from the initial state satisfies the temporal logic (TL) specification $\Phi$. Using these formal representations, we evaluate the videos.
@@
Text prompts are converted to temporal logic conversion using LLMs [chen2024nl2tltransformingnaturallanguages, cosler2023nl2specinteractivelytranslatingunstructured, Mendoza2024SYNTHTL, yang2023specification, sp_2025_CVPR] for subsequent analysis. We denote the LLM-based text-to-temporal logic (T2TL) function as $f_{\text{T2TL}}:T\to(\mathcal{P},\Phi)$ to decompose a text prompt $T$ into a TL specification $\Phi$ and a set of propositions $\mathcal{P}$. A prompt example is presented in the Appendix.
@@
Definition 1(Video Automaton Construction)
Given a generated video $\mathcal{V}$ and a set of atomic propositions $\mathcal{P}$, we construct a video automaton. For each proposition $p_{i}\in\mathcal{P}$ and frame $\mathcal{F}_{n}\in\mathcal{V}$, a VLM computes a semantic confidence score: $\mathcal{M}_{\text{VLM}}:p_{i}\times\mathcal{F}_{n}\to c_{i,n}$, where $c_{i,n}\in[0,1]$ represents the confidence of $p_{i}$ in $\mathcal{F}_{n}$. For each frame $n$, we define the confidence set as: $\mathbb{C}=\{c_{i,n}\mid p_{i}\in\mathcal{P},\mathcal{F}_{n}\in\mathcal{V}\ \}$. The video automaton $\mathcal{A}_{\mathcal{V}}$ is then generated by applying a function $\xi$ that processes the set of propositions $\mathcal{P}$ along with the confidence scores across all frames:
@@
Definition 2(Satisfaction Probability)
Given a video automaton $\mathcal{A}_{\mathcal{V}}$ and a temporal logic specification $\Phi$, the satisfaction probability $\mathbb{P}[\mathcal{A}_{\mathcal{V}}\models\Phi]$ is computed by verifying $\mathcal{A}_{\mathcal{V}}$ against $\Phi$ using STORM [storm, stormpy], which uses probabilistic computation tree logic (PCTL). This verification process is formalized as a probabilistic model checking function: